perform this using the linear differential equation of higher order in operator form: (D2 + 3D + 2) (e^-2x + 3x²)

1
Expert's answer
2021-11-30T16:39:32-0500

"\begin{aligned}
&\left(D^{2}+3 D+2\right)\left(e^{x}-2 x+3 x^{2}\right)\\
&=D^{2}\left(e^{x}-2 x+3 x^{2}\right)+3 D\left(e^{x}-2 x+3 n^{2}\right)
+2\left(e^{x}-2 x+3 x^{2}\right)\\
&= \frac{d^{2}}{d x^{2}}\left(e^{x}-2 x+3 x^{2}\right)+3 \frac{d}{d x}\left(e^{x}-2 x+3 x^{2}\right)
+2\left(e^{x}-2 x+3 x^{2}\right)\\
&=\frac{d}{d x}\left[\frac{d}{d x}\left(e^{x}-2 x+3 x^{2}\right)\right]+3\left(e^{x}-2+6 x\right)
+2\left(e^{x}-2 x+3 x^{2}\right)\\
&=\frac{d}{d x}\left[e^{x}-2+6 x\right]+3 e^{x}-6+18 x+2 e^{x}-4 x+6 x^{2}\\
&=e^{x}+6+3 e^{x}-6+18 x+2 e^{x}-4 x+6 x^{2}\\
&=6 e^{x}+14 x+6 x^{2} 
\end{aligned}"